A researcher classifies students according to their residence (urban or rural) and hours of paid work per week. He randomly selects one student. Residence 0 1–10 11–20 Over 20 Urban 0.06 0.20 0.15 0.06 Rural 0.12 0.18 0.20 0.03 a) If the student selected is from a rural region, what is the probability that he works between 1 and 10 hours a week? b) If the selected student works more than 20 hours a week, what is the probability that the student resides in an urban region?


1
Expert's answer
2021-07-06T18:19:51-0400

a)

A: student from rural region

"P(A)=0.12+0.18+0.20+0.03=0.53"

B: He works in 1-1 hours

"P(B)=0.18\\P(\frac{B}{A})=\frac{0.18}{0.53}=0.3396"


Approximately 0.34


b)

A: work more than 20 hours

"P(A)=0.061+0.03=0.09"

B: Besides in urban region

"P(B)=0 .06\\P(\frac{B}{A})=\frac{0.06}{0.09}=0.6666"

Approximately 0.067